### 12 Dec 2023 v1

# Pre-trained Universal Medical Image Transformer

Abstract
Self-supervised learning has emerged as a viable method to leverage the abundance of unlabeled medical imaging data, addressing the challenge of labeled data scarcity in medical image analysis. In particular, masked image modeling (MIM) with visual token reconstruction has shown promising results in the general computer vision (CV) domain and serves as a candidate for medical image analysis. However, the presence of heterogeneous 2D and 3D medical images often limits the volume and diversity of training data that can be effectively used for a single model structure. In this work, we propose a spatially adaptive convolution (SAC) module, which adaptively adjusts convolution parameters based on the voxel spacing of the input images. Employing this SAC module, we build a universal visual tokenizer and a universal Vision Transformer (ViT) capable of effectively processing a wide range of medical images with various imaging modalities and spatial properties. Moreover, in order to enhance the robustness of the visual tokenizer’s reconstruction objective for MIM, we suggest to generalize the discrete token output of the visual tokenizer to a probabilistic soft token. We show that the generalized soft token representation can be effectively integrated with the prior distribution regularization through a constructive interpretation. As a result, we pre-train a universal visual tokenizer followed by a universal ViT via visual token reconstruction on 55 public medical image datasets, comprising over 9 million 2D slices (including over 48,000 3D images). This represents the largest, most comprehensive, and diverse dataset for pre-training 3D medical image models to our knowledge. Experimental results on downstream medical image classification and segmentation tasks demonstrate the superior performance of our model and improved label efficiency. Our code is available at https://github.com/function2-llx/PUMIT.
Introduction
Over the last decade, medical image analysis, a field crucial for disease diagnosis and treatment planning, has notably benefited from the development of deep learning and shown success in various tasks (Litjens et al. 2017). Training these deep learning models requires a substantial amount of labeled data, while the availability of labeled medical data is often limited due to the expensive acquisition process. To address this issue, there has been a growing interest in self-supervised learning approaches, which learn representations from the large amounts of unlabeled data by designing self-supervised pretext tasks (Shurrab and Duwairi 2022). Despite the abundance of unlabeled data, the medical images are highly heterogeneous in terms of imaging modalities (CT, MRI, ultrasound, etc.) and, especially, spatial properties (2D and 3D spatial dimension, various voxel spacing and spatial shape), making it non-trivial to process all kinds of medical images with a unified model structure. The mainstream models are commonly designed to handle images of a single spatial property for a single model structure (Isensee et al. 2021). As a result, most of the previous works only leverage medical images data of similar spatial properties, limiting the amount and diversity of data for pre-training.
Recent advancements of the Vision Transformer (ViT) (Dosovitskiy et al. 2021) in the field of computer vision offer a promising solution to this challenge. The ViT converts the 2D or 3D spatial structure of an image into 1D patch sequences using a patch embedding layer and processes these sequences with transformer blocks. It has shown great potential in being used as a universal model to handle a wide range of medical images with diverse spatial properties. A recent work, UniMiSS (Xie et al. 2022a), has made initial attempts in this direction, leveraging both 2D and 3D unlabeled medical images for self-supervised pre-training. The authors designed a switchable patch embedding (SPE) module composed of both 2D and 3D patch embedding layers, which adaptively switches to either layer depending on the input dimension. While UniMiSS represents a significant step forward in breaking the “dimensionality barrier”, its patch embedding layer does not adaptively handle 3D medical images of various voxel spacings.
Following the success in the natural language processing (NLP) domain, masked image modeling (MIM) with ViT has emerged as a powerful self-supervised approach for learning vision representations. During MIM pre-training, an image is divided into non-overlapping patches (Dosovitskiy et al. 2021), with a subset being randomly masked out. The ViT model is tasked with predicting information related to the masked patches based on the visible ones. Currently, MIM pretext tasks are primarily categorized into three approaches: 1) pixel (voxel) value regression (Xie et al. 2022b), 2) visual token reconstruction with a pre-trained visual tokenizer (Bao et al. 2022), and 3) vision feature prediction (Wei et al. 2022). While pixel value regression is straightforward for MIM, it tends to focus on low-level information, given the inherent sparsity of image information (He et al. 2022; Fang et al. 2023b). Although the other two approaches of MIM can effectively capture high-level semantics, their application to medical image analysis is limited by the lack of universal models.
In this work, we aim to bridge the aforementioned gap by pre-training a universal visual tokenizer followed by a universal ViT with MIM via visual token reconstruction on a large volume of medical images. To achieve this, we introduce a spatially adaptive convolution (SAC) module, which adaptively adjusts convolution parameters based on the voxel spacing of the input image. Models built with the SAC module are capable of effectively processing a wide range of medical images with various imaging modalities and spatial properties. We also propose several improvements to the visual tokenizer to enhance the robustness of the reconstruction objective for MIM. A visual tokenizer is typically implemented using a vector-quantized variational autoencoder (VQ-VAE) (van den Oord, Vinyals, and Kavukcuoglu 2017). However, it can suffer from issues of severe codebook collapse and low codebook utilization with deterministic quantization, or the perturbations to the reconstruction process introduced by stochastic quantization during pre-training (Zhang et al. 2023). To address these issues, we generalize the discrete token to the soft token, which is represented as a categorical distribution over the codebook. We demonstrate the prior distribution regularization (Zhang et al. 2023) can be generalized to such soft token representation through a constructive interpretation. Furthermore, we extend this technique by maximizing the KL divergence of individual token distributions from the prior distribution, refining the sharpness of the learned token distributions. The contributions of our work are summarized as follows:
  • We propose the spatially adaptive convolution module that serves as a building block of the universal models for medical images. Consequently, we build such universal visual tokenizer and ViT with it for pre-training.
  • We generalize the discrete token in the visual tokenizer to the probabilistic soft token, mitigating codebook collapse without stochastic quantization. The token distributions are learned to be both diverse and sharpened through the extended prior distribution regularization.
  • We train the visual tokenizer followed by the ViT on 55 public medical image datasets, with over 9 million 2D slices (including over 48,000 3D images). This represents the largest, most comprehensive, and diverse dataset for pre-training 3D medical image models to our knowledge.
  • Our pre-trained models are fine-tuned on downstream medical image classification and segmentation tasks. Experimental results demonstrate the superior performance of our model and improved label efficiency.
Related Works
Self-supervised Learning on Medical Images
Self-supervised learning (SSL) aims to learn meaningful representations from unlabeled data by formulated pretext tasks. The application of SSL in medical image analysis can be mainly classified into the following three categories (Shurrab and Duwairi 2022). Contrastive SSL is a widely used technique that commonly models image similarity and dissimilarity through a family of data augmentations (Azizi et al. 2021; Sowrirajan et al. 2010; He et al. 2023). Predictive SSL aims to learn robust representations by formulating the pretext task as a classification or regression problem, such as geometric transformation prediction (Zhou et al. 2023), rotation angle prediction (Zhu et al. 2020; Tang et al. 2022), and context reconstruction (Zhou et al. 2021a; Chen et al. 2019). Generative SSL enables models to learn representations by either reconstructing input data or generating new examples from input distribution, commonly with autoencoders (Zhou et al. 2021b) and GANs (Liu et al. 2021). Recent works also explore leveraging auto-generated volumetric shapes for 3D medical image segmentation (Tadokoro, Yamada, and Kataoka 2023).
Masked Image Modeling
Inspired by the masked language modeling (MLM) in NLP (Devlin et al. 2019), masked image modeling (MIM) with transformer-based models has emerged as a viable SSL approach in the CV domain. Some representative works include SimMIM (Xie et al. 2022b) and MAE (He et al. 2022) with pixel value regression, BEiT family (Bao et al. 2022; Peng et al. 2022) and iBOT (Zhou et al. 2022) with visual token reconstruction, and MVP (Wei et al. 2022) and EVA series (Fang et al. 2023b, a) with vision feature prediction. Recent works also explore learning multimodal representations with MIM (Baevski et al. 2022, 2023) and investigate the scaling behaviors of MIM (Xie et al. 2023). However, there have been fewer attempts to apply MIM to medical image analysis (Chen et al. 2023), largely due to the significant differences between medical images and natural images.
Universal Models for Medical Images
A universal model capable of processing medical images with diverse properties under a unified model structure is crucial for leveraging a large amount of unlabeled data available for SSL. One straightforward approach to build such a universal model is to use a 2D model to process multiple 2D slices split from a 3D volume (Nguyen et al. 2023b, a). However, using a 2D model also requires additional mechanism to address the cross-plane contextual information inherent in 3D images. UniMiSS (Xie et al. 2022a) presents an alternative approach based on the ViT model, which converts the input 2D or 3D image into sequences of 1D image patches using a switchable patch embedding (SPE) layer. These sequences are then processed by transformer blocks. Additionally, the weight inflation technique (Zhang et al. 2022), which adapts pre-trained weights from 2D models to 3D models, also provides meaningful insights into the development of universal models for medical images.
Preliminary
In this section, we briefly review the concept of the visual tokenizer. Similar to a tokenizer in NLP, a visual tokenizer can convert an image $x$ into a $D\times H\times W$ grid of visual tokens, where each token belongs to a finite set $\mathcal{V}$, referred to as the visual codebook. Such a visual tokenizer can be implemented using a VQ-VAE (van den Oord, Vinyals, and Kavukcuoglu 2017), a variant of VAE (Kingma and Welling 2013) where the latent variables follow discrete categorical distributions over $\mathcal{V}$. Specifically, an encoder learns a posterior distribution $q(z\mid x)$ to output a $D\times H\times W$ grid of categorical token distributions over $\mathcal{V}$ given the image $x$. Token samples can be drawn from these categorical distributions in a differentiable manner using Gumbel-Softmax reparameterization (Maddison, Mnih, and Teh 2017; Jang, Gu, and Poole 2017) and straight-through estimators (Bengio, Léonard, and Courville 2013). Each token in $\mathcal{V}$ corresponds to a $d$-dimensional embedding vector from an embedding space $\bm{e}\in\mathbb{R}^{|\mathcal{V}|\times d}$. A decoder learns to reconstruct the original image given the embeddings of the drawn token samples. The prior distribution of the latent variable at each position is the uniform distribution over $\mathcal{V}$. Similarly to vanilla VAE, such visual tokenizer can be trained by maximizing the evidence lower bound (ELBO).
Methodology
Spatially Adaptive Convolution
Voxel spacing is an important spatial property of 3D medical images, representing the physical size of a voxel. Conventional image-processing networks, such as convolutional networks, operate on the voxel grid but disregard the voxel spacing information. To address this discrepancy, a common strategy is to resample the images in a dataset to a uniform spacing (Isensee et al. 2021). However, this approach is impractical for pre-training on a large volume of images with diverse voxel spacings.
Instead, we propose the adaptive adjustment of the convolution parameters based on the voxel spacing of the input. Typically, a 3D medical image consists of a stack of multiple slices, with voxel spacing defined by two degrees of freedom: the slice thickness $s_{\mathrm{slice}}$ and the isotropic in-plane spacing $s_{\mathrm{plane}}$. A 2D image can be viewed as a special case with only one slice and sufficiently large slice thickness. To explain how the SAC module works, we define the degree of anisotropy (DA) of a feature map to be $\mathrm{DA}=\max\left\{0,\left\lfloor\log_{2}\frac{{s_{\mathrm{slice}}}}{s_{\mathrm{plane}}}\right\rfloor\right\}$. The SAC module is essentially a 3D convolution or transposed convolution initialized with kernel sizes and strides that are isotropic across all dimensions, with the following three variants:
  • Downsampling: Both convolution kernel size and stride are $2^{k}$, where $k$ is a non-negative integer. The convolution weight along the depth dimension is reduced to the size of $2^{\max\{k-\mathrm{DA},0\}}$ through sum pooling, and the stride of the depth dimension is similarly adjusted to $2^{\max\{k-\mathrm{DA},0\}}$. The spacing of the output feature map is multiplied by the adjusted stride.
  • $3^{3}$ convolution: The convolution kernel size is 3 and stride is 1 on the depth dimension. If $\mathrm{DA}>0$, the convolution weight is reduced along the whole depth dimension to the size of 1 by sum pooling.
  • Upsampling: Similar to downsampling, both transposed convolution kernel size and stride are $2^{k}$, where $k$ is a non-negative integer, but the reduced kernel size and stride of the depth dimension is $2^{\min\{k,\mathrm{DA}_{0}-\mathrm{DA}\}}$, where $\mathrm{DA}_{0}$ is the DA of the input image. The spacing of the output feature map is divided by the adjusted stride.
We specifically employ sum pooling for weight reduction to maintain the mean and variance of the output, similar to the weight inflation technique used to adapt pre-trained weights from 2D models to 3D models (Zhang et al. 2022).
Universal Visual Tokenizer
A convolution-based universal visual tokenizer for medical images can be constructed with the proposed SAC module. The encoder of the tokenizer resembles a typical convolutional backbone, but with each $3^{3}$ convolution and downsampling convolution replaced by the corresponding variant of the SAC module. Given an input image with voxel spacing information, it outputs a feature map with $2^{-4}$ of the original resolution on the in-plane dimensions and $2^{-\max\{4-\mathrm{DA},0\}}$ of the original resolution on the depth dimension. The decoder is composed of a stack of the $3^{3}$ convolution and the upsampling variants of the SAC module. It gradually upsamples the quantized feature map output by the encoder, and finally outputs a reconstruction of the original image.
Universal ViT
The patch embedding layer of a ViT divides the input image into a sequence of non-overlapping patches, and maps each patch to an embedding vector through a linear projection. This process is equivalent to processing the input image with a downsampling convolution that has identical kernel size and stride, and flattening the spatial dimensions of the convolution output. Therefore, the downsampling convolution can be replaced by the downsampling variant of the SAC module to accommodate spatially heterogeneous input images. Furthermore, the transformer blocks following the patch embedding layer in ViT are inherently able to process sequences with variable lengths. As a result, a ViT enhanced with the proposed SAC module in the patch embedding layer becomes a universal backbone capable of handling a wide range of medical images with various imaging modalities and spatial properties.
Figure 1: Illustration of our proposed method for pre-training a universal visual tokenizer and a universal ViT for medical images. Our model can process a wide range of images with a unified model structure based on the proposed SAC module.
Visual Tokenizer Details
Existing pre-trained visual tokenizers can suffer from issues of severe codebook collapse, low codebook utilization, and perturbed reconstruction process introduced by stochastic quantization (Zhang et al. 2023). We propose several enhancements to make the visual tokenizer produce more robust tokenization output.
Soft Token Representation
Instead of sampling a discrete token from the categorical distribution, we directly represent a soft token with the distribution. The embedding of a soft token is calculated as the expectation of the token embedding. Specifically, let random variable $I$ be the index of a token in the codebook. The soft token embedding is given as:
$\mathrm{E}\left(\bm{e}_{I}\right)=\sum_{i=1}^{|\mathcal{V}|}\mathrm{P}(I=i)\bm{e}_{i}.$ (1)
This strategy could mitigate the codebook collapse issue, while still preserving a deterministic reconstruction process.
Extended Prior Distribution Regularization
During experiments, we find that using the soft token representation alone still results in low codebook utilization. The prior distribution regularization (Zhang et al. 2023) provides a possible direction of solution. It calculates a distribution $p_{\mathrm{post}}$ by averaging the one-hot distributions over the codebook of all sampled discrete tokens, and minimizes $D_{\mathrm{KL}}\left(p_{\mathrm{post}}\parallel p_{\mathrm{prior}}\right)$, where $p_{\mathrm{prior}}$ is a uniform distribution. This technique can be effortlessly generalized to the soft token representation by replacing the one-hot distributions with the general categorical distributions of tokens.
In fact, there exists a constructive interpretation when using this technique with the soft token representation. Suppose the encoder outputs a $D\times H\times W$ grid of token distributions, let $G_{D,H,W}$ denote the set of discrete coordinates of all cells within the grid. For each position $s\in G_{D,H,W}$ on the grid, we define a random variable $I_{s}$ denoting the index of the token in the codebook at $s$ and its distribution to be $q^{(s)}$. We define another random variable $S$ representing a random position on the grid that is uniformly distributed over $G_{D,H,W}$. It can be shown that the averaged distribution of all token distributions on the grid is exactly the distribution of $I_{S}$:
$\mathrm{P}(I_{S}=i)$ $=\sum_{s\in G_{D,H,W}}\mathrm{P}(I_{S}=i\mid S=s)\mathrm{P}(S=s)$ (2)
$=\frac{1}{DHW}\sum_{s\in G_{D,H,W}}\mathrm{P}(I_{s}=i).$ (3)
Intuitively, when the distribution of $I_{S}$ is close to the uniform distribution, individual tokens occur at a randomly sampled position with close probabilities. This benefits the diversity of the learned distributions.
On the other hand, to prevent all distributions collapsing to the uniform distribution, we introduce another objective maximizing $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ to increase the sharpness of the learned distributions. Since $p_{\mathrm{prior}}$ is a uniform distribution, it can be shown that for any distribution $p$, $D_{\mathrm{KL}}\left(p\parallel p_{\mathrm{prior}}\right)=-\mathrm{H}\left(p\right)+\ln\left\lvert\mathcal{V}\right\rvert$, where $\mathrm{H}\left(p\right)=-\sum_{i=1}^{|\mathcal{V}|}p_{i}\ln p_{i}$. Therefore, the extended prior distribution regularization can be implemented as minimizing the following objective:
$L_{\mathrm{reg}}=-\lambda_{1}\mathrm{H}\left(\mathrm{E}\left(q^{(S)}\right)\right)+\lambda_{2}\mathrm{E}\left(\mathrm{H}\left(q^{(S)}\right)\right),$ (4)
where $\lambda_{1},\lambda_{2}$ are positive hyper-parameters.
Optimization Objective
Besides the basic L1 reconstruction loss $L_{\mathrm{rec}}$, the training objective of our visual tokenizer includes a perceptual loss $L_{\mathrm{percep}}$ (Zhang et al. 2018) and a discriminative loss $L_{\mathrm{disc}}$ by adversarily training a patch-based distriminator $P$ following VQGAN, and the aforementioned extended prior distribution regularization $L_{\mathrm{reg}}$. During our experiments, we find that using the gradient-based adaptive weight for $L_{\mathrm{disc}}$ proposed in VQGAN results in failure of convergence. Instead, we set the weight for $L_{\mathrm{disc}}$ as a fixed hyper-parameter $\lambda$, and include $L_{\mathrm{disc}}$ in the total objective only when the loss function value of the discriminator $L_{P}$ is smaller than a given threshold $t_{\mathrm{disc}}$. In conclusion, the final optimization objective is minimizing:
$L_{\mathrm{VT}}=L_{\mathrm{rec}}+L_{\mathrm{percep}}+[L_{P}<t_{\mathrm{disc}}]L_{\mathrm{disc}}+L_{\mathrm{reg}},$ (5)
where the []delimited-[][$\cdot$][ ⋅ ] operator is Iverson bracket. The overall structure of our universal visual tokenizer is illustrated in the upper part of Figure 1.
Transformer Architecture
Our ViT model employs a plain ViT macro architecture that is composed of a patch embedding layer and a stack of transformer blocks (Dosovitskiy et al. 2021). The micro architecture of the transformer block follows EVA-02 (Fang et al. 2023a), which incorporates several advancements in language models for visual representation learning, including the 2D rotary position embedding (RoPE) (Su et al. 2021). In short, for each $d$-dimensional query or key vector $\bm{a}=(a_{0},a_{1},\dots,a_{d-1})$ in an attention block, 1D RoPE rotates the pair $(a_{2i},a_{2i+1})$ on the 2D plane by a radian of $\omega_{i}t$, where $t$ is the position index, and it is commonly chosen that $\omega_{i}=b^{-\frac{2i}{d}}$ with a base $b$. For a given position on a 2D image coordinated with $(t_{x},t_{y})$, 2D RoPE extends 1D RoPE by rotating the first half of the vector according to $t_{x}$ and the second half according to $t_{y}$.
We further extend 2D RoPE to work with 3D medical images for any 3D coordinate $(t_{x},t_{y},t_{z})$ by adding an additional rotation radian. Specifically, let $\theta_{i}$ be the rotation radian given by the 2D RoPE at $(t_{x},t_{y})$ for the pair $(a_{2i},a_{2i+1})$, the total rotation radian is $\theta_{i}+\omega_{z,i}t_{z}$, where $\omega_{z,i}=b_{z}^{-\frac{2i+1}{d}}$ with a base $b_{z}$. Further details of RoPE can be found in the appendix.
Figure 2: Reconstruction results of our visual tokenizer on unseen images. The images are taken from different body parts with diverse imaging modalities and spatial properties. For each pair of images, the original image is on the left and the reconstructed image is on the right. The first row: (left) chest X-ray image from CheXpert (Irvin et al. 2019); (middle) fundus photograph from GAMMA; (right) breast ultrasound image from BUSI (Al-Dhabyani et al. 2020). The second row: (left): lung CT image from LIDC-IDRI (Armato III et al. 2011); (middle): abdomen CT image from FLARE 2022; (right): prostate MRI image from PROSTATE-MRI (Choyke et al. 2016). We clarify that the results of 3D images in the second row are obtained by processing the volume as a whole with the visual tokenizer and selecting a slice from the reconstructed volume. More examples are available in the appendix.
MIM Pre-training
We pre-train the ViT via visual token reconstruction with our pre-trained visual tokenizer, which is similar to BEiT (Bao et al. 2022) but the visual tokens are in soft representations. The overall method is illustrated in Figure 1. For a given 3D image $x$, it is converted into a grid of token distributions $\left\{q^{(s)}\mid s\in G_{D,H,W}\right\}$ by the pre-trained visual tokenizer. On the other hand, the patch embedding layer of ViT also divides $x$ into corresponding $D\times H\times W$ patches. We randomly mask out 55% of the patches and replace their embeddings with a learnable embedding $\bm{e}_{\mathrm{mask}}$. Transformer blocks then process the corrupted patch embeddings and output a $D\times H\times W$ grid of hidden vectors. These hidden vectors are projected to the categorical distributions $\left\{\tilde{q}^{(s)}\mid s\in G_{D,H,W}\right\}$ over $\mathcal{V}$ through a multi-layer perceptron (MLP) with a single hidden layer. The pre-training objective is minimizing the cross entropy of the reconstructed distribution relative to the distribution output by the visual tokenizer:
$L_{\mathrm{MIM}}=\frac{1}{\left\lvert\mathcal{M}\right\rvert}\sum_{s\in\mathcal{M}}\sum_{i=1}^{\left\lvert\mathcal{V}\right\rvert}q^{(s)}_{i}\ln\tilde{q}^{(s)}_{i},$ (6)
where $\mathcal{M}\subset G_{D,H,W}$ is the set of positions of all masked patches.
We intentionally choose a higher masking ratio for 3D medical images compared to previous works on 2D images (Bao et al. 2022; Fang et al. 2023a). In addition to patches within the same plane, patches from other planes may also provide useful information for reconstruction. A higher masking ratio helps balance the difficulty of the pre-training task.
Experiments
Pre-training
Dataset
We collect 55 public medical image datasets for both visual tokenizer and ViT pre-training, including 15 datasets of 2D images and 40 datasets of 3D images. The collected data include a wide range of image modalities (CT, MRI, X-ray, ultrasound and fundus photography) and body parts (abdomen, brain, breast, cervix, chest, fetal, fundus, head-neck, heart, pelvis, prostate and spine), with diverse spatial properties. We compare the amount and imaging modalities of data for pre-training with related works in Table 1. We pre-process the data to a unified format, including moving the depth dimension of an image to the first spatial dimension. During pre-training, we only use raw images in a self-supervised manner, and any annotations or labels from the datasets are not utilized. For each modality in each dataset, we randomly exclude one image from training data for validation. The details of all datasets for pre-training and the way of pre-processing are provided in the appendix.
Work #3D #2D Modality
Swin UNETR 5,050 0 CT
SMIT 3,643 0 CT
Nguyen et al. 3,694 108,948 CT, XR
UniMiSS 5,022 108,948 CT, XR
Ours 48,344 348,991
CT, MRI,
XR, US, FP
Table 1: The amount and imaging modalities of data used for pre-training in related and our works. Note: #3D and #2D: the number of 3D and 2D images, respectively; XR: X-ray; US: ultrasound; FP: fundus photography.
Implementation Details
For visual tokenizer, we set codebook size $|\mathcal{V}|=1024$. During pre-training, we randomly crop a $\max\{64/2^{\mathrm{DA}},1\}\times 128\times 128$ patch from an image given its DA. We adopt a batch sampling strategy that all images within a batch are selected to have the same DA, ensuring that they have the same cropped size and thus can be processed efficiently by the network in parallel. We train the visual tokenizer for 600K steps with the batch size of 48 and the peak learning rate of $10^{-4}$. For ViT pre-training, we randomly crop a $\max\{80/2^{\mathrm{DA}},1\}\times 160\times 160$ patch from an image, and use the same batch sampling strategy as the visual tokenizer training. We initialize ViT with the EVA-02-B pre-trained weights and train it for 300K steps with the batch size of 84 and the peak learning rate of $2\times 10^{-5}$. All pre-trained models are trained with AdamW optimizer (Loshchilov and Hutter 2019) and cosine learning rate scheduler with a linear warm-up. Our models are implemented with PyTorch (Paszke et al. 2019) and MONAI (Cardoso et al. 2022) and trained with NVIDIA A100 GPUs.
Results
We demonstrate the reconstruction results of our visual tokenizer on unseen validation images in Figure 2. Our universal visual tokenizer could process a wide range of medical images with heterogeneous imaging modalities and spatial properties with a unified model structure. As shown in Figure 2, images from diverse sources can be consistently reconstructed from the learned soft token representations with rich details preserved. Moreover, we conduct the following analyses to evaluate the diversity and sharpness of the learned distributions. First, following Zhang et al., we plot the codebook utilization on validation images in Figure 3. The usage percentages of all tokens are at the same order of magnitude, indicating a balanced utilization of the codebook. Second, we calculate the average value of $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ on validation images as 2.791, while $\mathrm{H}\left(p_{\mathrm{prior}}\right)=\ln|\mathcal{V}|\approx 6.931$. This result indicates that the learned distributions preserve sharpness to some extent.
Figure 3: Token utilization histogram. Every eight tokens with consecutive indexes are merged into one rectangle for better visualization in limited page width.
Classification Experiments
Dataset
MedMNIST v2 (Yang, Shi, and Ni 2021; Yang et al. 2023) is a large-scale benchmark for medical image classification on standardized MNIST-like 2D and 3D images with diverse modalities, dataset scales and tasks. It consists of 12 datasets with 708,069 2D images and 6 datasets with 9,998 3D images. The size of each image is either $28\times 28$ (2D) or $28\times 28\times 28$ (3D). We experiment on the 3D portion of MedMNIST v2, namely the Organ (Bilic, Christ et al. 2019; Xu, Zhou et al. 2019), Nodule (Armato III et al. 2011), Fracture (Jin et al. 2020a), Adrenal, Vessel (Yang et al. 2020) and Synapse datasets.
Implementation Details
Like common ViT-based models, our model performs classification tasks by feeding the hidden state of the [CLS] token into a fully-connected layer. We conduct classification experiments by fine-tuning from our and several other self-supervised pre-trained models on each dataset. Following the official split of training/validation/test sets, we train all models on the training sets for 100 epochs. The model weights from the epoch with the highest AUC score on the validation set are selected for comparison on the test set. The details on experimental settings are given in the appendix.
Methods Organ Nodule Fracture Adrenal Vessel Synapse Average
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet-50+3D 99.4 88.3 87.5 84.7 72.5 49.4 82.8 74.5 90.7 91.8 85.1 79.5 86.3 78.0
ResNet-50+ACS 99.4 88.9 88.6 84.1 75.0 51.7 82.8 75.8 91.2 85.8 71.9 70.9 84.8 76.2
auto-sklearn 97.7 81.4 91.4 87.4 62.8 45.3 82.8 80.2 91.0 91.5 63.1 73.0 81.5 76.5
UniMiSS 99.8 94.9 89.9 84.2 70.4 50.8 87.0 81.2 93.8 94.0 85.8 81.8 87.8 81.2
SMIT 99.5 91.1 88.4 85.2 63.1 49.6 84.8 82.2 89.7 91.1 70.8 73.3 82.7 78.8
EVA-02-B 99.2 88.0 91.0 85.8 75.2 56.7 86.7 79.5 95.8 95.6 72.1 75.3 86.7 80.2
Ours 99.8 95.6 94.3 88.4 77.6 55.4 87.8 81.2 97.3 95.6 96.4 91.5 92.2 84.6
Table 2: Classification results on the test sets of the 3D image datasets from MedMNIST v2. The results of ResNet-50 and auto-sklearn are taken from the original paper (Yang et al. 2023). Both AUC and ACC values are shown in percentages.
Results
Table 2 compares the classification performance of our model with other models using area under the ROC curve (AUC) and accuracy (ACC) metrics. Our model consistently outperforms all other models across all datasets in terms of AUC, and achieves the highest ACC on 4 out of 6 datasets. On average, our model improves AUC by at least 4.4% and accuracy by at least 3.4% across all datasets.
Segmentation Experiments
Dataset
We use the following two datasets to evaluate the segmentation performance. 1) The Beyond the Cranial Vault (BTCV) dataset, which contains 30 CT scans with 13 classes of regions annotated. We follow the train/validation split of SMIT (Jiang et al. 2022) to use 21 scans for training and 9 scans for validation. 2) The MRI subset of Combined Healthy Abdominal Organ Segmentation (CHAOS) dataset, which contains 40 pairs of T1-DUAL MRI images and 40 T2-SPIR MRI images with 4 classes of regions annotated. The segmentation labels of 20 pairs of T1-DUAL images and 20 T2-SPIR images are available for training, while the labels for remaining cases are hidden for online testing.
Implementation Details
For both tasks, we resample the images to the median spacing among the training data. Our pre-trained ViT is used to initialize the backbone, which extracts feature maps with the fixed in-plane scale of $\frac{1}{16}$. To build hierarchical feature maps for the segmentation head, following the implementation of EVA-02 (Fang et al. 2023a), we process the extracted feature maps with transposed convolutions to obtain feature maps with the in-plane scales of $\frac{1}{8}$ and $\frac{1}{4}$, and a max-pooling layer to obtain a feature map with the in-plane scale of $\frac{1}{32}$. Subsequently, a plain U-Net decoder processes the hierarchical feature maps to obtain the segmentation output. All segmentation models for both tasks are trained for 50K steps and the final weights are used for evaluation. Dice coefficient, average symmetric surface distance (ASSD), and Hausdorff distance (HD) are used to evaluate results for both tasks. The details on experimental settings and evaluation metrics are given in the appendix.
Results
Method BTCV (validation) CHAOS (hidden test)
Dice ASSD HD Dice ASSD HD
TFS 84.86 1.59 18.90 89.96 2.18 25.39
Swin UNETR 81.06 3.20 92.33 85.51 6.66 68.02
UniMiSS 85.17 1.01 17.78 90.49 1.91 24.94
EVA-02-B 86.51 0.86 16.91 91.48 1.71 20.69
Ours 87.31 0.74 13.12 92.09 1.54 19.85
Table 3: Segmentation results on the validation set of BTCV and hidden test set of CHAOS. “TFS” indicates the model with a plane ViT backbone trained from scratch. The values of Dice are shown in percentages.
Table 3 compares the segmentation performance of our model with other models on the BTCV validation set and CHAOS hidden test set. Our model outperforms all other models in terms of all 3 metrics for both tasks. Notably, although EVA-02-B is pre-trained on natural images, fine-tuning from it could already outperform Swin UNETR (Tang et al. 2022) and UniMiSS that are pre-trained on medical images. We hypothesize that this is due to the insufficient medical image data used for pre-training.
$r$ Dice ASSD HD
TFS Ours $\Delta$ TFS Ours $\Delta$ TFS Ours $\Delta$
0.2 70.98 77.84 6.85 4.96 2.39 2.57 51.13 25.56 25.57
0.4 75.11 81.28 6.17 4.67 2.41 2.26 43.85 21.71 22.14
0.6 80.73 84.87 4.14 2.74 1.23 1.51 31.70 18.03 13.67
0.8 84.73 86.96 2.23 1.61 0.78 0.83 23.78 13.66 10.12
1.0 84.86 87.31 2.45 1.59 0.74 0.86 18.90 13.12 5.78
Table 4: Segmentation results on the BTCV validation set comparing models trained from scratch (TFS) with those fine-tuned from our pre-trained weights, using different ratios $r$ of training data. The columns of “$\Delta$” indicate the amount of improvement.
Ablation Study
To investigate how our model improves the label efficiency of the downstream tasks, we compare the model trained from scratch and fine-tuned from our model with different ratios of training data in Table 4. Fine-tuning from our pre-trained model consistently outperforms training from scratch with a large margin across all metrics, and the improvements are more significant when less data are available. Moreover, our pre-trained model could effectively mitigate the requirement for labeled data. As shown in Table 4, even when fine-tuning with 60% of the training data, our pre-trained model still achieves better results than a model trained from scratch using the full dataset.
Conclusion
In this work, we introduce the SAC module, which serves as the building block for universal models for medical images. We are able to leverage 55 datasets for pre-training, which comprises over 9 million 2D slices (including over 48,000 3D images), representing the most extensive pre-training effort for 3D medical image models, to the best of our knowledge. Additionally, we develop a universal visual tokenizer for medical images, offering as a robust reconstruction objective for MIM. Our pre-trained model is demonstrated with superior downstream performance and label efficiency.
Appendix
Details on RoPE
From 1D to 2D RoPE
Positional embeddings play a crucial role in transformer models (Vaswani et al. 2017). RoPE (Su et al. 2021) is a kind of position embedding based on rotation transformation. We extend the 2D rotary position embedding (RoPE) (Su et al. 2021) to work with 3D medical images.
For a $d$-dimensional query or key vector $\bm{a}=(a_{0},a_{1},\dots,a_{d-1})$ in an attention block, the vanilla RoPE, primarily designed for sequence data, rotates each pair $(a_{2i},a_{2i+1})(i=0,1,\cdots,\frac{d}{2}-1)$ on the 2D plane by a radian of $\omega_{i}t$, where $t$ is the position index and it is commonly chosen that $\omega_{i}=b^{-\frac{2i}{d}}$ with a base $b$. This can be represented by matrix multiplication $\mathcal{R}(t)\bm{a}$, the transformation matrix is defined as
$\mathcal{R}(t)=\mathrm{diag}\left(T(\omega_{0}t),\dots,T(\omega_{\frac{d}{2}-1}t)\right),$ (7)
where $T(\theta)$ is defined as the 2D rotation matrix that rotates a radian of $\theta$:
$T(\theta)=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}.$ (8)
When handling 2D image data, for a given position on a 2D image coordinated with $(t_{x},t_{y})$, 2D RoPE extends 1D RoPE by rotating the first half of the vector $\bm{a}$ according to $t_{x}$ and the second half according to $t_{y}$. The corresponding transformation matrix $\mathcal{R}(t_{x},t_{y})$ is given by:
$\mathcal{R}(t_{x},t_{y})=\begin{bmatrix}\mathcal{R}(t_{x})&\\ &\mathcal{R}(t_{y})\end{bmatrix}.$ (9)
Both vanilla 1D RoPE and extended 2D RoPE preserves the following properties. Take 2D RoPE for example:
  • Relativity. For two positions $(t_{x}^{(1)},t_{y}^{(1)})$ and $(t_{x}^{(2)},t_{y}^{(2)})$, it can be easily verified that $\mathcal{R}^{\mathsf{T}}(t_{x}^{(1)},t_{y}^{(1)})\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})=\mathcal{R}(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$. Consequently, for two vectors $\bm{a}$ and $\bm{b}$, the result of their inner product after applying RoPE is a quantity dependent on the relative position $(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$:
    $\left(\mathcal{R}(t_{x}^{(1)},t_{y}^{(1)})\bm{a}\right)^{\mathsf{T}}\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})\bm{b}$ (10)
    $=$ $\bm{a}^{\mathsf{T}}\mathcal{R}^{\mathsf{T}}(t_{x}^{(1)},t_{y}^{(1)})\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})\bm{b}$ (11)
    $=$ $\bm{a}^{\mathsf{T}}\mathcal{R}(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})\bm{b}.$ (12)
  • Reversibility. $\mathcal{R}(t_{x}^{(1)},t_{y}^{(1)})=\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})$ if and only if $(t_{x}^{(1)},t_{y}^{(1)})=(t_{x}^{(2)},t_{y}^{(2)})$, as long as the value of each $\omega_{i}$ are appropriately chosen.
Additive 3D RoPE
We aim to extend RoPE to 3D images to handle the 3D coordinate $(t_{x},t_{y},t_{z})$, while maintaining the aforementioned properties of relativity and reversibility. Furthermore, to better utilize the pre-trained weights of 2D models, we also seek to ensure that when the relative position on the $z$-axis is 0, the results are equal to the 2D counterparts. Specifically, the transformation matrix at $(t_{x},t_{y},t_{z})$ is given as:
$\mathcal{R}(t_{x},t_{y},t_{z})=\begin{bmatrix}\mathcal{R}_{x,z}(t_{x},t_{z})&\\ &\mathcal{R}_{y,z}(t_{y},t_{z})\end{bmatrix},$ (13)
where
$\forall a\in\{x,y\},\mathcal{R}_{a,z}(t_{a},t_{z})=$ (14)
$\mathrm{diag}\left(T(\omega_{a,0}t_{a}+\omega_{z,0}t_{z}),\dots,T(\omega_{a,\frac{d}{2}-1}t_{a}+\omega_{z,\frac{d}{2}-1}t_{z})\right).$
In our experiments, we set $\omega_{x,i}=b_{x}^{-\frac{2i}{d}},\omega_{y,i}=b_{y}^{-\frac{2i}{d}},\omega_{z,i}=b_{z}^{-\frac{2i+1}{d}}$, where $b_{x}=b_{y}=10000,b_{z}=2333$. We show with our specifically chosen $\omega_{x,i},\omega_{y,i},\omega_{z,i}$, such approach maintains the properties of relativity and reversibility.
Relativity. We only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies relativity. For each diagonal entry, we have:
$T^{\mathsf{T}}(\omega_{x,i}t_{x}^{(1)}+\omega_{z,i}t_{z}^{(1)})T(\omega_{x,i}t_{x}^{(2)}+\omega_{z,i}t_{z}^{(2)})$ (15)
$=$ $T(\omega_{x,i}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z,i}(t_{z}^{(2)}-t_{z}^{(1)})).$ (16)
Therefore, $\mathcal{R}^{\mathsf{T}}_{x,z}(t_{x}^{(1)},t_{z}^{(1)})\mathcal{R}_{x,z}(t_{x}^{(2)},t_{z}^{(2)})=\mathcal{R}_{x,z}(t_{x}^{(2)}-t_{x}^{(1)},t_{z}^{(2)}-t_{z}^{(1)})$.
Reversibility. Similarly, we only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies reversibility. Note that the values we specifically choose for $\omega_{x,i},\omega_{y,i},\omega_{z,i}$ are algebraic numbers, the following theorem helps proves the result.
Theorem 0.1.
If both $\omega_{x}$ and $\omega_{z}$ are non-zero algebraic numbers, and $\frac{\omega_{x}}{\omega_{z}}$ is an irrational number, then for any $(t_{x}^{(1)},t_{z}^{(1)}),(t_{x}^{(2)},t_{z}^{(2)})\in\mathbb{Z}^{2}$ where $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$, it holds that $\mathcal{R}_{x,z}(t_{x}^{(1)},t_{z}^{(1)})\neq\mathcal{R}_{x,z}(t_{x}^{(2)},t_{z}^{(2)})$
Proof.
The proposition is equivalent to:
$\omega_{x}t_{x}^{(1)}+\omega_{z}t_{z}^{(1)}\not\equiv\omega_{x}t_{x}^{(2)}+\omega_{z}t_{z}^{(2)}\pmod{2\pi}.$ (17)
Proof by contradiction. Assume there exists $(t_{x}^{(1)},t_{z}^{(1)}),(t_{x}^{(2)},t_{z}^{(2)})\in\mathbb{Z}^{2}$ and $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$, $k\in\mathbb{Z}$ such that
$2k\pi=\omega_{x}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z}(t_{z}^{(2)}-t_{z}^{(1)}).$ (18)
If $k=0$, then $t_{x}^{(2)}\neq t_{x}^{(1)}$ (otherwise, from $0=\omega_{x}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z}(t_{z}^{(2)}-t_{z}^{(1)})$, we would also have $t_{z}^{(2)}=t_{z}^{(1)}$, which contradicts $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$). Therefore, from (18) we have:
$\frac{\omega_{x}}{\omega_{z}}=-\frac{t_{z}^{(2)}-t_{z}^{(1)}}{t_{x}^{(2)}-t_{x}^{(1)}}\in\mathbb{Q},$ (19)
which contradicts the condition that $\frac{\omega_{x}}{\omega_{z}}$ is irrational.
If $k\neq 0$, since the product and sum of algebraic numbers are still algebraic, from (18), we have
$\pi=\dfrac{\omega_{x}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z}(t_{z}^{(2)}-t_{z}^{(1)})}{2k}$ is also algebraic. However, this contradicts the well-known fact that $\pi$ is transcendental. ∎
Experiment
To validate the effectiveness of the proposed Additive 3D RoPE, we computed the rotation radians corresponding to all positions on images of a typical size. Note that the $x$ and $y$ dimensions are symmetric in Additive 3D RoPE, therefore, in our experiments, we set the length of the $y$ dimension to 1, and only focus on the $x$ and $z$ dimensions. Accordingly, for each $10\leq i\leq d/4-$ 1 where $d$ is the dimension of the query or key vector $\bm{a}$), we examined the distribution of these radians in terms of $\bmod 2\pi$ representation.
For image size to be validated, we set it as $8\times 16\times 1(z\times x\times y)$. Notice that that when the z-coordinate $t_{z}=0$, the rotation radian becomes $\omega_{x,i}t_{x}+\omega_{z,i}t_{z}=\omega_{x,i}t_{x}$, which is equivalent to 2D RoPE. Therefore, we also choose 2D image size of $1\times 16\times 1(z\times x\times y)$ for comparison with 2D RoPE.
For each $10\leq i\leq d/4-$ 1, we consider the results of rotating radians of all positions modulo $2\pi$ to obtain $\Theta_{i}$. Subsequently, we calculate the differences between all adjacent elements after sorting $\Theta_{i}$, resulting in $\Delta_{i}$. Finally, we calculate the minimum value among these differences as $\min\Delta_{i}$. The formal definitions of $\Theta_{i},\Delta_{i}$ are as follows:
$\Theta_{i}$ =$=\{(\omega_{x,i}t_{x}+\omega_{z,i}t_{z})\bmod 2\pi\mid 0\leq t_{x}<x,0\leq t_{z}<z\}$
=$=\{\theta_{1},\theta_{2},\cdots,\theta_{xz}\}\quad(0\leq\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{xz}<2\pi)$
$\Delta_{i}$ =$=\{\theta_{i+1}-\theta_{i}\mid 1\leq i\leq xz-1\}\cup\{\theta_{0}+2\pi-\theta_{xz}\}$
In addition, we also attempt to use the Euclidean distance between the positions after rotating $(0,1)\in\mathbb{R}^{2}$ by radians of $\alpha$ and $\beta$ as an indicator to quantify the ”difference” between these two rotation radians, as follows:
[…]
$\mid 1\leq i\leq xz\},$
where $\theta_{xz+1}$ is defined as $\theta_{1}$.
Since there is a considerable number of $\theta_{i}$ values, the distribution of them on the unit circle $x^{2}+y^{2}=1$ is relatively dense. In this scenario, $\sqrt{(\cos\theta_{i+1}-\cos\theta_{i})^{2}+(\sin\theta_{i+1}-\sin\theta_{i})^{2}}\approx(\theta_{i+1}-\theta_{i})\bmod 2\pi$, implying that the arc length corresponding to this central angle will be approximately equal to the chord length. Consequently, the results obtained from these two evaluation metrics will be quite close. For simplicity, only the differences of adjacent radians are presented below.
Figure 4: Minimum intervals between rotation radians for different $i$. $i$ is the index of dimensions $d$ of query or key vector $\bm{a}$.
It can be observed that when using Additive 3D RoPE on 3D images, the minimum intervals between rotation radians are approximately in the range of $10^{-3}$ to $10^{-5}$, which preserves good numerical precision with 32bit floating point numbers. As for the comparison with 2D RoPE, the average result of $\frac{\min\Delta_{i}^{(2D)}}{\min\Delta_{i}^{(3D)}}$ for all $i$ is 26.30, which benefits 3D medical images processing. By setting $b_{z}<b_{x}$, which implies $\omega_{z,i}>\omega_{x,i}$, Additive 3D RoPE gains anisotropy, allowing it to carry cross-plane positional information. The effect of this choice on the distribution of rotation radians is that rotation radians for positions with the same $z$ coordinate tend to cluster together, while positions with different $z$ coordinates in the entire image give rise to distinct clusters of rotation radians. Thus, the difference between our results and the ideal ratio of 16.00 for both cases with uniform distribution reflects the anisotropic information.
As shown in Figure 4, we can also observe the relationship between $b_{z}$ and $\min\Delta_{i}$: if $b_{z}$ is large, indicating that $\omega_{z,i}$ is small, all rotation radians are densely distributed, resulting in a smaller $\min\Delta_{i}$; whereas if $b_{z}$ is small, although $\min\Delta_{i}$ is better, larger $\omega_{z,i}$ can lead to rotation radians for positions that are far apart become too close in terms $\bmod 2\pi$, which contradicts the relativity pursued by RoPE. In conclusion, we finally set $b_{z}=2333$ in our work.
Configuration UniMiSS SMIT EVA-02-B Ours
initial learning rate 5e-8 / 1e-7 / 5e-7 / 1e-6 / 5e-6 / 1e-5 / 5e-5 / 1e-4 / 5e-4
lr decay factor 0.1
training epochs 100
lr schedule lr step at epoch 50, 75
batch size 32
optimizer AdamW (Loshchilov and Hutter 2019)
weight decay 1e-2
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
fine-tuning input size $96\times 96\times 96$ $96\times 96\times 96$ $28\times 224\times 224$ $160\times 160\times 160$
pre-training input size $16\times 96\times 96$ $96\times 96\times 96$ $224\times 224$ $80\times 160\times 160$
Table 5: Classification fine-tuning settings
Pre-training Data and Pre-processing
From each dataset, we exclude any corrupted data or those lacking consistent spacing information. During pre-processing, for each image, we determine its depth dimension, i.e., the dimension along which the image is scanned, and move the depth dimension to the first spatial dimension. The depth dimension is determined based on the following priorities: 1) adhering to official data description if available; 2) extracting from image metadata if available; 3) using the dimension with the most anisotropic spacing compared to other dimensions. We crop the image to the region with intensities above zero for RGB and gray scale images, and above the 0.5 percentile for others. If the shorter in-plane size is greater than 512, we resize the image to have a shorter in-plane size of 512 with area interpolation. The full list of datasets and the amount of data in each dataset used for pre-training are presented in Table LABEL:table:datasets. The number of slices for 3D images is counted along the depth dimension. In total, 48,344 3D images and 348,991 2D images, amounting to 9,096,684 slices, are used for pre-training.
More Reconstruction Results
We show more visual tokenizer reconstruction results in Figure 5. Our visual tokenizer is able to reconstruct various anatomical structure with diverse imaging modalities and spatial properties.
Classification Experiments
Experimental Settings
Our experimental settings on MedMNIST v2 largely follow the official baselines (Yang et al. 2023), including the regularization for the Adrenal and Vessel datasets. For fair comparisons, we search for optimal learning rate when fine-tuning each model on each dataset. To utilize the pre-trained weights of models, we adopt the following adjustments. First, for each model, each input image is interpolated into the size that is compatible with the input size used by the model during pre-training. Second, the 3D images of MedMNIST v2 have only one channel. For each model that is pre-trained on images with three channels (such as RGB images), we duplicate each input image along the channel dimension into three channels. The detailed recipe is provided in Table 5.
Segmentation Experiments
Experimental Settings
For experiments on BTCV, a $48\times 192\times 192$ patch from each input image during training for all models but Swin UNETR. The model structure of Swin UNETR only supports input size that is divisible by 32, and we thus follow the official implementation of Swin UNETR [^1] https://github.com/Project-MONAI/research-contributions/tree/main/SwinUNETR/BTCV to use a crop size of $96\times 96\times 96$ specifically for it. Following Swin UNETR, we use a soft tissue window of $[-175,250]$ for clipping the CT intensities, then rescale intensities to [-1, 1]. For experiments on CHAOS, the crop size during training is $32\times 192\times 192$ for all models. For each MRI image, we normalize it by subtracting the mean of intensities and dividing standard deviation of intensities. For both tasks, all models are trained for 50K steps and segmentation results are obtained by sliding window inference with a overlap ratio of 0.75. The remained hyper-parameters settings for each model follow the corresponding official implementation.
Evaluation Metrics
Given a prediction set of points $\tilde{V}$ and a reference set $V$, we evaluate the segmentation results using the Dice coefficient, average symmetric surface distance (ASSD), and Hausdorff distance (HD) as follows:
$\mathrm{Dice}(\tilde{V},V)=\frac{2\lvert\tilde{V}\cap V\rvert}{\lvert V\rvert+\lvert\tilde{V}\rvert},$ (20)
[…]
$\mathrm{HD}(\tilde{V},V)=\max\left\{\max_{s\in\partial V}d(s,\partial\tilde{V}),\max_{s\in\partial\tilde{V}}d(s,\partial V)\right\},$ (22)
where $\partial V$ is the boundary of $V$, and
$d(s,V)=\min_{t\in V}d(s,t),$ (23)
where $d(s,t)$ is the Euclidean distance between points $s$ and $t$. More specific details of the evaluation metrics can be found in the official documentation of the CHAOS challenge [^2] https://github.com/emrekavur/CHAOS-evaluation/blob/master/CHAOS˙Metrics˙and˙Evaluation˙Method.pdf. For results on BTCV validation set, we use the MONAI’s implementation for the metrics to evaluate them. For results on CHAOS test set, we submit them to the challenge [^3] https://chaos.grand-challenge.org and report the evaluation results returned by the website.
Figure 5: More visual tokenizer reconstruction results on unseen images. The first row: (left) abdomen CT image MSD-Pancreas; (middle) fetal ultrasound image from HC18; (right) head-and-neck MRI image from HaN-Seg. The second row: (left): heart MRI image from MSD-Heart; (middle): abdomen CT image from FLARE 2022; (right): pelvix X-ray image from PelviXNet. The third row: (left): lung image from STOIC; (middle): spine CT image from RSNA-2022-CSFD; (right): brain MRI image from crossMoDA2022. The forth row: (from left to right) T1-weighted, T2-weighted, and T2-weighted FLAIR brain MRI images from BraTS-MEN. The fifth row: (from left to right) fundus photograph from iChallenge-ADAM, iChallenge-REFUGE2, and RFMiD, respectively.
[…]
datasets.tex